# InstructLR: A Scalable Approach to Create Instruction Dataset for Under-Resourced Languages

### v1

@@ Abstract
Mamadou K. KEITA1, Sebastien Diarra2, Christopher Homan1, Seydou Diallo3, 1Rochester Institute of Technology, 2RobotsMali 3MALIBA-AI
@@ 1 Introduction
Large language models (LLMs) are proficient in many tasks, with recent models sometimes outperforming humans, depending on the language. They tend to perform substantially worse on low-resource languages (LRLs), such as those spoken across Africa and other regions, than on higher-resource languages. This performance gap is evidently due to the limited representation of these languages in pre-training and fine-tuning datasets. Although LLMs such as GPT-4 (openai2024gpt4technicalreport) and Gemini (geminiteam2024geminifamilyhighlycapable) have made progress in multilingual capabilities, many LRLs remain poorly, if at all, supported.
Existing approaches to address this gap also face major limitations. Machine translation (MT) of fine-tuning datasets from higher-resourced languages into LRLs often produces unnatural text that fails to capture language-specific nuances (zhu-etal-2024-multilingual). Synthetic data generation frequently results in hallucinated content and a lack of cultural awareness (guo2024generativeaisyntheticdata). The relatively high cost of creating human-annotated instruction data for LRLs worsens the situation.
@@ 2.1 Seed Instruction
To ensure the final dataset is comprehensive and useful for training models, InstructLR starts by curating a diverse set of topics. We draw inspiration from established multi-task benchmarks like MMLU (hendrycks2021measuringmassivemultitasklanguage) because they provide a structured framework of knowledge domains and reasoning skills. Our selection process targets a balanced distribution across a wide range of areas. These include STEM fields (e.g., Physics, Mathematics, Computer Science), humanities (e.g., History, Law, Philosophy), and social sciences. The goal is to create a dataset that supports not only knowledge recall but also the development of complex reasoning abilities.
@@
After gathering the topic list, seed instructions are generated in a high-resource language. This approach is a necessary adaptation of the self-instruct method (wang-etal-2023-self-instruct) for the LRL context. The standard self-instruct loop is technically infeasible here, as it requires a teacher model with strong generative capabilities in the target language to create novel instructions—a prerequisite that current models do not meet for languages like Zarma. Our method circumvents this by using the LLM for the task it can perform well (ideation in French). The choice of the high-resource language depends on its presence in the region where the target LRL is used—e.g., French-speaking countries will use French.
@@ 3 Dataset Creation and Analysis
To demonstrate the effectiveness of InstructLR for generating instruction datasets, we report on our use of it to create a dataset in Zarma, a West African language spoken by over six million people (keita-etal-2024-feriji).
@@ 3.1 Seed Instruction Creation
For this stage, we selected 20 topics— listed with descriptions in Table 10—and proceeded with instruction generation. Since Zarma coexists with French in everyday usage (keita-etal-2024-feriji), we chose French as the primary language for generating seed instructions, and a suitable model for French: the Mistral 7b model (jiang2023mistral7b). We then generated French instructions per topic and equally split across the topics ($\approx$ 5% per topic).
@@ 3.2 Draft Generation
Once we had the curated set of French seed instructions and their associated topics, we moved on to generating the first drafts of instruction-response pairs in Zarma  [^2] All $50,000$ instructions were processed, and a snapshot of the outputs is shown in Table 11.. To achieve this, we tested several models—Gemini 2.5 Pro, GPT 4.o, and Llama 3.3 (grattafiori2024llama3herdmodels)—to determine which one demonstrated a relatively acceptable understanding of Zarma.
@@ 3.3 Quality Assessment
Knowledge base construction: Our RAG checker used a knowledge base of 3,000 clean sentences from the Feriji dataset (keita-etal-2024-feriji), 20 Zarma grammar rules each followed by examples, and bilingual glossaries, all encoded with a faiss dense index (douze2025faisslibrary). This knowledge base enabled the system to contextualize and evaluate drafts with high precision.
@@ 4 Experiments
We evaluate six open-source models across different parameter scales: Gemma-3-270M, Gemma-3-1B, Gemma-3-4B gemmateam2025gemma3technicalreport, Llama-3.1-8B grattafiori2024llama3herdmodels, Mistral-7B-Instruct-v0.3, and Phi-4 abdin2024phi4technicalreport. For each language, we split our 50k datasets into 49,000 training pairs and 1,000 held-out test pairs for evaluation.
For the baselines, We compare against two baselines; Zero-Shot Baseline: Each base model evaluated on test sets without fine-tuning. MT-Seed Baseline: To isolate the effect of our generation pipeline, we create a controlled comparison using direct MT of our French seed instructions. We fine-tune Llama-3.1-8B (our best model across all the languages experimented before the MT one)on datasets created by translating the same 50,000 French seed instructions using MADLAD-400 (kudugunta2023madlad400)—because MADLAD is the only known model (untill this date) that supports all the three languages of this experiment. This approach avoids confusion caused by culture-specific instructions in existing datasets such as Alpaca (alpaca).
We use unsloth (unsloth) with QLoRA (dettmers2023qloraefficientfinetuningquantized) for efficient fine-tuning. Training parameters include: learning rate 2e-5, 3 epochs, with CoT responses included as supervised targets. We ensure no overlap between training and test sets.
@@
Lang InstructLR vs. InstructLR Baseline Ties
@@
Table 3: Results of the human preferences experiment. The human evaluation and the MT-Seed were carried out with our best-performing model (Llama-3.1-8B with InstructLR).
Table 2 presents results on held-out test sets using BLEU (papineni-etal-2002-bleu), ROUGE-L (lin-2004-rouge), and METEOR (banerjee-lavie-2005-meteor) metrics. Zero-shot performance demonstrates limitations of current LLMs for these languages, with scores near zero across all models—which confirms that Zarma, Bambara, and Fulfulde are minimally or not covered by the models used for the trainings.
@@ Appendix B Related Work
Instruction tuning aligns LLMs with user needs by fine-tuning on task instruction data (ma2025buildinginstructiontuningdatasetshumanwritten). Benchmarks—like FLAN, T0, etc—provide instruction datasets for LLMs to be trained on (wei2022finetunedlanguagemodelszeroshot; sanh2022multitaskpromptedtrainingenables; wang2024mmluprorobustchallengingmultitask; hendrycks2021measuringmassivemultitasklanguage; wang2020supergluestickierbenchmarkgeneralpurpose; wang2019gluemultitaskbenchmarkanalysis). However, these advances are centered on higher-resource languages—leaving LRLs with marginal coverage. This is particularly true for many African languages, due to the lack of task-specific data and the affordability of creating data. Recent work addresses this gap through multilingual instruction tuning. muennighoff-etal-2023-crosslingual showed that fine-tuning a multilingual model on English tasks can enable zero-shot instruction-following in other languages present only in pre-training data. Moreover, adding a small portion of multilingual data during fine-tuning yields further improvements on the target-language tasks (muennighoff-etal-2023-crosslingual). Nevertheless, “severely” LRLs—particularly African languages—still lag behind, as the current benchmarks cover only relatively better-represented languages—such as Hausa or Swahili.
Several works provide instruction data specifically for African languages. For instance, Masakhane has produced datasets for tasks such as machine translation (MT) or named entity recognition (e.g., MasakhaNER supports 10 African languages (adelani2021masakhanernamedentityrecognition)). AfriInstruct integrates translation data (FLORES, MAFAND-MT for 16 languages), topic classification and summarization data (XL-Sum, etc), sentiment corpora (AfriSenti and NollySenti), and Masakhane benchmarks (NER, POS tagging) into a unified training set (uemura-etal-2024-afriinstruct). Yet, these are limited to a few African languages—not even half of the total languages present in the region. Our work addresses the need for scale-appropriate tools for building instruction datasets for LRL.
@@
Due to the lack of human-written instruction data in most LRLs, a popular alternative is synthetic instruction generation. The self-instruct framework proposed by wang-etal-2023-self-instruct demonstrated that one can create an instruction dataset by prompting a language model with a handful of seed tasks to produce new instruction–response pairs. Following this, researchers have explored extending self-instruction to other languages. For example, chen-etal-2024-monolingual translates the Alpaca English instructions into eight languages to compare multilingual vs. monolingual instruction tuning, and finds that even machine-translated instructions can provide cross-lingual benefits.
Also, it is important to mention the recent trend of using LLMs as annotators to reduce the cost of creating LRL data. For instance, alhanai2024bridginggapenhancingllm leverage GPT-4o to automate parts of their quality assessment process by having the model score generated text on metrics such as fluency and factual consistency.
However, purely synthetic data approaches are not fully reliable in terms of quality. Model-generated instructions may contain errors, non-fluent phrasing, or cultural inappropriateness in the target LRL. Recent work highlights the need for careful control of LLM-synthesized data using strategies like rewriting the generated instructions or having multiple LLMs chat with each other to stimulate feedback dialog (ma2025buildinginstructiontuningdatasetshumanwritten). Despite these solutions, this limitation still remains, and proves the need of human-in-the-loop approaches within these processes.
@@ C.1 System Architecture
The RAG checker integrates two primary components: a retrieval module and a generation/assessment module. The retrieval module uses a knowledge base comprising 3,000 clean Zarma sentences from the Feriji dataset (keita-etal-2024-feriji), 20 Zarma grammar rules with examples, and bilingual glossaries. These resources were encoded with a FAISS dense index (douze2025faisslibrary) for efficient semantic retrieval.
@@ C.2 Evaluation Protocols
To evaluate the RAG checker, we designed a controlled test set of 300 Zarma sentences. The test set comprised 200 sentences with injected grammatical errors, created by prompting the DeepSeek v3 (deepseekai2025deepseekv3technicalreport) LLM to break specific Zarma grammar rules, and 100 unaltered sentences that served as a gold standard for measuring false-positive rates. Each sentence was processed through the RAG analyzer, and the system’s assessments and corrections were compared with the gold references.

### v2

@@ Abstract
Mamadou K. KEITA1, Sebastien Diarra2, Christopher Homan1, Seydou Diallo3 1Rochester Institute of Technology, 2RobotsMali 3MALIBA-AI
@@ 1 Introduction
Large language models (LLMs) are proficient in many tasks, with recent models sometimes outperforming humans, depending on the language. They tend to perform substantially worse on low-resource languages (LRLs), such as those spoken across Africa and other regions, than on higher-resource languages. This performance gap is evidently due to the limited representation of these languages in pre-training and fine-tuning datasets. Although LLMs such as GPT-4 (OpenAI et al., 2024) and Gemini (Team et al., 2024) have made progress in multilingual capabilities, many LRLs remain poorly, if at all, supported.
Existing approaches to address this gap also face major limitations. Machine translation (MT) of fine-tuning datasets from higher-resourced languages into LRLs often produces unnatural text that fails to capture language-specific nuances (Zhu et al., 2024). Synthetic data generation frequently results in hallucinated content and a lack of cultural awareness (Guo and Chen, 2024). The relatively high cost of creating human-annotated instruction data for LRLs worsens the situation.
@@ 2.1 Seed Instruction
To ensure the final dataset is comprehensive and useful for training models, InstructLR starts by curating a diverse set of topics. We draw inspiration from established multi-task benchmarks like MMLU (Hendrycks et al., 2021) because they provide a structured framework of knowledge domains and reasoning skills. Our selection process targets a balanced distribution across a wide range of areas. These include STEM fields (e.g., Physics, Mathematics, Computer Science), humanities (e.g., History, Law, Philosophy), and social sciences. The goal is to create a dataset that supports not only knowledge recall but also the development of complex reasoning abilities.
@@
After gathering the topic list, seed instructions are generated in a high-resource language. This approach is a necessary adaptation of the self-instruct method (Wang et al., 2023) for the LRL context. The standard self-instruct loop is technically infeasible here, as it requires a teacher model with strong generative capabilities in the target language to create novel instructions—a prerequisite that current models do not meet for languages like Zarma. Our method circumvents this by using the LLM for the task it can perform well (ideation in French). The choice of the high-resource language depends on its presence in the region where the target LRL is used—e.g., French-speaking countries will use French.
@@ 3 Dataset Creation and Analysis
To demonstrate the effectiveness of InstructLR for generating instruction datasets, we report on our use of it to create a dataset in Zarma, a West African language spoken by over six million people (Keita et al., 2024).
@@ 3.1 Seed Instruction Creation
For this stage, we selected 20 topics— listed with descriptions in Table 10—and proceeded with instruction generation. Since Zarma coexists with French in everyday usage (Keita et al., 2024), we chose French as the primary language for generating seed instructions, and a suitable model for French: the Mistral 7b model (Jiang et al., 2023). We then generated French instructions per topic and equally split across the topics ($\approx$ 5% per topic).
@@ 3.2 Draft Generation
Once we had the curated set of French seed instructions and their associated topics, we moved on to generating the first drafts of instruction-response pairs in Zarma  [^2] All $50,000$ instructions were processed, and a snapshot of the outputs is shown in Table 11.. To achieve this, we tested several models—Gemini 2.5 Pro, GPT 4.o, and Llama 3.3 (Grattafiori et al., 2024)—to determine which one demonstrated a relatively acceptable understanding of Zarma.
@@ 3.3 Quality Assessment
Knowledge base construction: Our RAG checker used a knowledge base of 3,000 clean sentences from the Feriji dataset (Keita et al., 2024), 20 Zarma grammar rules each followed by examples, and bilingual glossaries, all encoded with a faiss dense index (Douze et al., 2025). This knowledge base enabled the system to contextualize and evaluate drafts with high precision.
@@ 4 Experiments
We evaluate six open-source models across different parameter scales: Gemma-3-270M, Gemma-3-1B, Gemma-3-4B Team et al. (2025), Llama-3.1-8B Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3, and Phi-4 Abdin et al. (2024). For each language, we split our 50k datasets into 49,000 training pairs and 1,000 held-out test pairs for evaluation.
For the baselines, We compare against two baselines; Zero-Shot Baseline: Each base model evaluated on test sets without fine-tuning. MT-Seed Baseline: To isolate the effect of our generation pipeline, we create a controlled comparison using direct MT of our French seed instructions. We fine-tune Llama-3.1-8B (our best model across all the languages experimented before the MT one)on datasets created by translating the same 50,000 French seed instructions using MADLAD-400 (Kudugunta et al., 2023)—because MADLAD is the only known model (untill this date) that supports all the three languages of this experiment. This approach avoids confusion caused by culture-specific instructions in existing datasets such as Alpaca (Taori et al., 2023).
We use unsloth (Daniel Han and team, 2023) with QLoRA (Dettmers et al., 2023) for efficient fine-tuning. Training parameters include: learning rate 2e-5, 3 epochs, with CoT responses included as supervised targets. We ensure no overlap between training and test sets.
@@
Lang Comparison InstructLR Baseline Ties
@@
Table 3: Results of the human preferences experiment. Human evaluation and MT-Seed were carried out with our best-performing model (Llama-3.1-8B with InstructLR).
Table 2 presents results on held-out test sets using BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) metrics. Zero-shot performance demonstrates limitations of current LLMs for these languages, with scores near zero across all models—which confirms that Zarma, Bambara, and Fulfulde are minimally or not covered by the models used for the trainings.
@@ Appendix B Related Work
Instruction tuning aligns LLMs with user needs by fine-tuning on task instruction data (Ma et al., 2025). Benchmarks—like FLAN, T0, etc—provide instruction datasets for LLMs to be trained on (Wei et al., 2022; Sanh et al., 2022; Wang et al., 2024; Hendrycks et al., 2021; Wang et al., 2020, 2019). However, these advances are centered on higher-resource languages—leaving LRLs with marginal coverage. This is particularly true for many African languages, due to the lack of task-specific data and the affordability of creating data. Recent work addresses this gap through multilingual instruction tuning. Muennighoff et al. (2023) showed that fine-tuning a multilingual model on English tasks can enable zero-shot instruction-following in other languages present only in pre-training data. Moreover, adding a small portion of multilingual data during fine-tuning yields further improvements on the target-language tasks (Muennighoff et al., 2023). Nevertheless, “severely” LRLs—particularly African languages—still lag behind, as the current benchmarks cover only relatively better-represented languages—such as Hausa or Swahili.
Several works provide instruction data specifically for African languages. For instance, Masakhane has produced datasets for tasks such as machine translation (MT) or named entity recognition (e.g., MasakhaNER supports 10 African languages (Adelani et al., 2021)). AfriInstruct integrates translation data (FLORES, MAFAND-MT for 16 languages), topic classification and summarization data (XL-Sum, etc), sentiment corpora (AfriSenti and NollySenti), and Masakhane benchmarks (NER, POS tagging) into a unified training set (Uemura et al., 2024). Yet, these are limited to a few African languages—not even half of the total languages present in the region. Our work addresses the need for scale-appropriate tools for building instruction datasets for LRL.
@@
Due to the lack of human-written instruction data in most LRLs, a popular alternative is synthetic instruction generation. The self-instruct framework proposed by Wang et al. (2023) demonstrated that one can create an instruction dataset by prompting a language model with a handful of seed tasks to produce new instruction–response pairs. Following this, researchers have explored extending self-instruction to other languages. For example, Chen et al. (2024) translates the Alpaca English instructions into eight languages to compare multilingual vs. monolingual instruction tuning, and finds that even machine-translated instructions can provide cross-lingual benefits.
Also, it is important to mention the recent trend of using LLMs as annotators to reduce the cost of creating LRL data. For instance, Alhanai et al. (2024) leverage GPT-4o to automate parts of their quality assessment process by having the model score generated text on metrics such as fluency and factual consistency.
However, purely synthetic data approaches are not fully reliable in terms of quality. Model-generated instructions may contain errors, non-fluent phrasing, or cultural inappropriateness in the target LRL. Recent work highlights the need for careful control of LLM-synthesized data using strategies like rewriting the generated instructions or having multiple LLMs chat with each other to stimulate feedback dialog (Ma et al., 2025). Despite these solutions, this limitation still remains, and proves the need of human-in-the-loop approaches within these processes.
@@ C.1 System Architecture
The RAG checker integrates two primary components: a retrieval module and a generation/assessment module. The retrieval module uses a knowledge base comprising 3,000 clean Zarma sentences from the Feriji dataset (Keita et al., 2024), 20 Zarma grammar rules with examples, and bilingual glossaries. These resources were encoded with a FAISS dense index (Douze et al., 2025) for efficient semantic retrieval.
@@ C.2 Evaluation Protocols
To evaluate the RAG checker, we designed a controlled test set of 300 Zarma sentences. The test set comprised 200 sentences with injected grammatical errors, created by prompting the DeepSeek v3 (DeepSeek-AI et al., 2025) LLM to break specific Zarma grammar rules, and 100 unaltered sentences that served as a gold standard for measuring false-positive rates. Each sentence was processed through the RAG analyzer, and the system’s assessments and corrections were compared with the gold references.